### v1

# Quantization-Aware Neuromorphic Architecture for Efficient Skin Disease Classification on Resource-Constrained Devices

Abstract
Accurate and efficient skin lesion classification on edge devices is critical for accessible dermatological care but remains challenging due to computational, energy, and privacy constraints. We introduce QANA, a novel quantization-aware neuromorphic architecture for incremental skin lesion classification on resource-limited hardware. QANA effectively integrates ghost modules, efficient channel attention, and squeeze-and-excitation blocks for robust feature representation with low-latency and energy-efficient inference. Its quantization-aware head and spike-compatible transformations enable seamless conversion to spiking neural networks (SNNs) and deployment on neuromorphic platforms. Evaluation on the large-scale HAM10000 benchmark and a real-world clinical dataset shows that QANA achieves 91.6% Top-1 accuracy and 82.4% macro F1 on HAM10000, and 90.8% Top-1 accuracy and 81.7% macro F1 on the clinical dataset, consistently outperforming leading CNN-to-SNN models under fair comparison. Deployed on BrainChip Akida hardware, QANA achieves 1.5 ms inference latency and 1.7 mJ energy per image, reducing inference latency and energy use by over 94.6%/98.6% compared to GPU-based CNNs, and exceeding the performance of advanced CNN-to-SNN conversion methods. These results demonstrate the effectiveness of QANA for accurate, real-time, and privacy-sensitive medical analysis in edge environments.
Index Terms:
Resource-Constrained Devices, Edge Computing, Neuromorphic Computing, Low Latency, Energy Efficiency.
I Introduction
Skin diseases present significant diagnostic challenges, particularly for clinicians without specialized training, leading to frequent misdiagnosis of conditions such as melanoma and Merkel cell carcinoma [1, 2]. Although deep learning-based diagnostic systems have shown promising performance [3], most approaches rely on centralized training and inference, requiring sensitive patient data to be transferred to cloud servers [4]. This raises data security risks [5] and is constrained by strict privacy regulations such as HIPAA [6] and GDPR [7]. Furthermore, to support dermatological care in home and remote settings lacking conventional healthcare infrastructure[8], it is essential to develop models that enable on-device training and inference.
However, deploying deep learning models on edge devices faces several challenges. Conventional CNNs require significant computational and energy resources, increasing device complexity and limiting portability [9]. In addition, effective training typically depends on large labeled datasets, which are often unavailable for rare skin diseases [10, 11]. Frequent model updates further introduce computational overhead and privacy concerns, while hardware constraints restrict model complexity and inference accuracy on edge platforms [2].
Figure 1: Detailed architecture of our end-to-end framework for quantization-aware neuromorphic skin lesion classification: (1) data preprocessing (quality filtering, augmentation, and SMOTE-based oversampling); (2) a novel quantization-aware network for feature extraction and spike-compatible transformation; (3) CNN-to-SNN conversion with operator mapping and temporal spike encoding; and (4) SNN deployment with on-chip optimization for real-time and energy-efficient inference on edge hardware.
Recently, Spiking Neural Networks (SNNs) and neuromorphic computing platforms have emerged as promising alternatives for overcoming deployment challenges in dermatological diagnosis [12]. Unlike conventional CNNs, SNNs use discrete spike events for information encoding and transmission, resulting in sparse and event-driven computation that greatly reduces power consumption [13]. This spike-based encoding is well-suited for scenarios with limited labeled data, as the temporal dynamics of spikes facilitate effective learning from fewer examples [14]. Neuromorphic processors, such as BrainChip’s Akida [15], IBM’s TrueNorth [16], and Intel’s Loihi [17], natively support SNNs and enable on-chip incremental learning. Incremental learning enables these systems to adapt efficiently to new patient data without full model retraining, aligning well with clinical practices that regularly acquire new diagnostic cases [18]. Additionally, these neuromorphic devices possess compact physical footprints and substantially lower energy requirements compared to conventional GPUs [19], making them viable candidates for edge-based deployment in portable diagnostic instruments.
The prevalent approach for neuromorphic deployment is to first train conventional CNN architectures, such as ResNet or DenseNet, on large datasets, and then convert them to equivalent SNN models for edge inference [20]. While this conversion aims to combine the accuracy of CNNs with the efficiency of SNNs [21], several practical limitations persist. Key CNN components, such as Batch Norm, Flatten, and Global Average Pooling, cannot be directly mapped to spike-based neuron units, complicating and often degrading the conversion process [22]. In addition, converted SNNs frequently experience significant accuracy loss on small, imbalanced medical datasets due to quantization effects and limited capacity for capturing subtle pathological features [23].
In this work, we present an end-to-end pipeline for efficient and incremental skin lesion classification on neuromorphic hardware. As illustrated in Fig. 1, our approach consists of four main stages: (1) data preprocessing, including image quality screening, augmentation, and SMOTE-based class balancing; (2) construction of a quantization-aware neural network with stacked Ghost modules, channel attention, and a spike-compatible output head; (3) automated conversion of the trained network to a spiking neural network (SNN) using the Akida MetaTF toolkit; and (4) deployment and optimization on the BrainChip Akida neuromorphic platform for low-latency, energy-efficient inference. The core QANA architecture is specifically designed for hardware compatibility, enabling direct and lossless translation to spike-based event-driven computation. This pipeline directly addresses the major challenges of edge deployment, including computational constraints, class imbalance, limited data, and the need for incremental model updates.
Experimental results on both the HAM10000 public benchmark and a real-world clinical dataset demonstrate the practical advantages of our approach. QANA achieves 91.6% Top-1 accuracy and 82.4% macro F1 on HAM10000, and 90.8% / 81.7% on the clinical dataset. On Akida hardware, the system delivers 1.5 ms inference latency and 1.7 mJ energy consumption per image, surpassing state-of-the-art CNN-to-SNN conversion baselines in both accuracy and efficiency. These findings confirm the effectiveness of QANA for resource-constrained, real-time medical image analysis.
II Related Work
Prior studies on neuromorphic hardware and Spiking Neural Networks (SNNs) have highlighted their suitability for edge inference and continual learning in constrained environments [14]. Platforms such as BrainChip’s Akida [15], IBM’s TrueNorth [16], and Intel’s Loihi [17] demonstrate ultra-low-power operation and support on-chip incremental learning, enabling model updates without full retraining. These features are advantageous for medical imaging tasks on edge devices. However, existing works primarily focus on general benchmarks like MNIST [24] or CIFAR [25] and rarely address domain-specific challenges in dermatology [26]. Furthermore, in medical image contexts, edge Spiking Neural learning methods often suffer from catastrophic forgetting and fail to preserve performance on previously learned classes when new samples are introduced [27].
CNN-to-SNN conversion has become a common strategy to leverage mature CNN architectures—ResNet, DenseNet—for deployment on neuromorphic processors [28]. Conversion toolkits, including Akida’s CNN2SNN [29], deliver compatibility by quantizing weights and replacing activation functions with spike-based equivalents. While some studies report preserved accuracy on standard datasets [30], performance significantly degrades on small or imbalanced medical datasets due to batching overfitting, quantization noise, and inability to capture rare lesion features [15]. Limitations in mapping CNN operations—such as batch normalization, global average pooling, and multi-bit activations—to spiking neurons further compound accuracy loss during conversion [31].
Figure 2: Detailed architecture of our QANA, which performs iterative feature extraction using stacked Ghost modules, ECA, and residual blocks, followed by spike-compatible transformation with batch normalization, ReLU activation, and Squeeze-and-Excitation (SE) block. The output is then quantized and projected to class logits for SNN deployment.
In parallel, conventional CNN-based methods for skin lesion classification, such as those using DenseNet-121 [32] or Inception-v4 [33], show high performance on benchmark datasets like ISIC. Despite their effectiveness, these models necessitate continuous retraining on cloud infrastructure to address domain shifts, new lesion types, or demographic variations [18]. Such retraining incurs latency, high computational cost, and data privacy concerns [34]. Moreover, general CNN models often fail to generalize to underrepresented conditions [35], such as rare tumors or images from diverse skin tones, highlighting the need for adaptive and privacy-preserving on-device solutions. These limitations motivate the development of neuromorphic frameworks that maintain diagnostic performance while supporting incremental learning and efficient operation in resource-constrained settings.
III Methodology
This section details the complete pipeline for the development and deployment of the proposed neuromorphic skin lesion classification system. As shown in Fig. 1, the pipeline encompasses data preprocessing, the design of a quantization-aware neural network architecture compatible with spiking inference, conversion to an event-driven SNN format, and hardware-level deployment on the Akida platform.
III-A Data Preprocessing
The pipeline includes three main stages: image preprocessing to ensure quality and compatibility with the Akida hardware, data augmentation to enhance dataset diversity and robustness, and SMOTE to address severe class imbalance among lesion categories.
III-A1 Image Preprocessing
All dermatoscopic images are first screened for quality. Corrupted, low-resolution, or artifact-contaminated images are excluded using automated checks based on image metadata and pixel statistics, followed by manual review when ambiguity remains. For compatibility with the Akida neuromorphic platform, all accepted images are resized to $64\times 64$ pixels using bilinear interpolation, which is the maximum supported resolution under current on-chip memory constraints. Pixel intensities are normalized channel-wise to the range $[0,1]$ to ensure consistent input distribution. No denoising or inpainting operations are performed, as quality control steps eliminate samples with significant artifacts.
III-A2 Data Augmentation
To mitigate overfitting and increase the diversity of the training dataset, each image in the training split undergoes stochastic augmentation with the following operations: brightness adjustment (random factor sampled uniformly from $[0.7,1.3]$), contrast modification ($[0.8,1.2]$), random horizontal and vertical flipping (each with probability 0.5), hue shift (random shift in $[-0.08,0.08]$ in HSV space), and saturation variation ($[0.85,1.15]$). Each augmentation is applied independently to each sample with the specified probability. All augmentation parameters are determined by fixed random seeds for reproducibility. These transformations simulate real-world imaging variation while preserving lesion semantics.
III-A3 Synthetic Minority Oversampling (SMOTE)
To address class imbalance, especially in rare lesion categories, the Synthetic Minority Oversampling Technique (SMOTE) [36] is applied to the training data. Given a minority class dataset $\mathcal{S}=\{x_{1},x_{2},...,x_{n}\}$ in $\mathbb{R}^{d}$, for each $x_{i}\in\mathcal{S}$, the $k$ nearest neighbors $\mathcal{N}_{i}=\{x_{i_{1}},...,x_{i_{k}}\}$ are identified in feature space via Euclidean distance. For each synthetic sample, a random neighbor $x_{i_{j}}$ is selected, and a new sample is generated by:
$x_{\text{new}}=x_{i}+\lambda\cdot(x_{i_{j}}-x_{i}),\quad\lambda\sim\mathcal{U}(0,1)$ (1)
where $\lambda$ is independently sampled for each feature dimension. Repeating this process for all $x_{i}$ and multiple neighbors results in the augmented set:
$\mathcal{S}_{\text{aug}}=\mathcal{S}\cup\left\{x_{i}+\lambda^{(m)}\cdot(x_{i_{j_{m}}}-x_{i})\mid x_{i}\in\mathcal{S},\,1\leq m\leq M\right\}$ (2)
where $M$ is the number of synthetic samples generated per original sample. All images are maintained at $64\times 64$ resolution to conform with Akida’s memory requirements throughout oversampling. SMOTE is applied only to the training set to preserve test set integrity and eliminate information leakage.
III-B Quantization-Aware Network Architecture
The model (as shown in Fig. 2) comprises a hierarchical cascade of Ghost-based multi-scale feature extraction blocks, integrated channel attention mechanisms, and a quantization-aware transformation stage that produces SNN-ready outputs. This architecture balances computational efficiency and discriminative power, facilitating direct conversion and robust inference on the Akida neuromorphic platform.
III-B1 Iterative Feature Extraction and Downsampling
Feature extraction begins with a stack of Ghost blocks [37], designed to maximize representational diversity while minimizing computational cost. As shown in Fig. 3, each Ghost block receives a feature tensor $F^{(l-1)}$ (for $l=1,\ldots,4$) and applies both base and ghost convolutions as:
$F_{\text{ghost}}^{(l)}=\text{Concat}\Big{(}$ $\underbrace{\mathcal{F}_{1\times 1}^{(l)}\big{(}F^{(l-1)}\big{)}}_{\text{base: pointwise conv}},\,\underbrace{\mathcal{F}_{3\times 3}^{(l)}\big{(}F^{(l-1)}\big{)}\odot\mathbf{M}^{(l)}}_{\text{ghost: depthwise conv with filter mask}}\Big{)}$ (3)
Here, $\mathcal{F}_{k\times k}^{(l)}$ is a separable convolution of kernel size $k\times k$, and $\mathbf{M}^{(l)}$ is a binary mask selecting learnable ghost filters. This combination generates both local and extended receptive fields, enabling richer features with fewer FLOPs.
After concatenation, the tensor is processed by batch normalization, quantization-bounded activation, and dropout:
$F_{\text{drop}}^{(l)}=\text{Dropout}\Big{(}\min\big{(}6,\,\max(0,\,\gamma^{(l)}\cdot\text{BN}(F_{\text{ghost}}^{(l)})+\beta^{(l)})\big{)}\Big{)}$ (4)
This quantization-aware step is crucial for downstream SNN compatibility and robust generalization.
Figure 3: Schematic of the Ghost module. The input feature map is first processed by a lightweight convolution to extract a reduced set of primary features with channel size $\mu C$, where $C$ is the target output dimensionality and $\mu\in(0,1)$ is a tunable ratio. Subsequently, inexpensive operations are applied to the primary features to generate additional ghost features of size $(1-\mu)C$. These are concatenated along the channel axis to form the final output of size $C$.
To capture cross-channel correlations, each block integrates a Spatially-Aware Efficient Channel Attention (SA-ECA) mechanism, inspired by Efficient Channel Attention (ECA) mechanism [38]. As shown in Fig. 4, Instead of computing global channel statistics via global average pooling and 1D convolution, we adopt a lightweight depthwise convolution followed by pointwise channel-wise scaling. This enables efficient modeling of spatial-channel dependencies with minimal overhead and neuromorphic compatibility:
$\widetilde{F}^{(l)}=\sigma\left(\mathbf{W}_{1\times 1}^{(l)}*\text{BN}\left(\text{DWConv}_{k\times k}\left(F^{(l)}\right)\right)\right)\odot F^{(l)}$ (5)
where $\text{DWConv}_{k\times k}$ and $\mathbf{W}_{1\times 1}^{(l)}$ jointly form the attention mechanism in our SA-ECA block, and $\sigma$ is the sigmoid activation used to generate the attention mask. This lightweight attention preserves channel expressiveness with minimal overhead.
To ensure stable deep stacking, we employ residual skip connections and spatial downsampling:
$F^{(l)}=\text{MaxPool2D}\left(\boldsymbol{\alpha}^{(l)}\odot F_{\text{drop}}^{(l)}+\mathbf{P}^{(l)}F^{(l-1)}\right)$ (6)
where $\mathbf{P}^{(l)}$ projects the previous block’s output for dimension alignment if needed. This design preserves gradient flow, reduces vanishing/exploding risk, and aggregates multi-scale context, all critical for reliable feature extraction in small-data regimes. After four such blocks, the spatial size is reduced to $4\times 4$.
Figure 4: Illustration of the Spatially-Aware ECA (SA-ECA) block. A depthwise convolution is first applied to extract channel-wise statistics, followed by a lightweight 1D convolution to model local channel dependencies. The resulting attention weights are used to rescale the input feature channels, enhancing discriminative information with minimal computational overhead.
Figure 5: Illustration of the Squeeze-and-Excitation (SE) block. The input feature map undergoes global pooling, followed by two fully connected layers with ReLU and sigmoid activations to compute channel-wise weights. The original feature map is then rescaled by these weights, enabling adaptive recalibration of channel responses.
III-B2 Spike-Compatible Feature Transformation
The output of the previous stage, $F^{(4)}$, is passed to a spike-compatible transformation module engineered for direct quantization and SNN integration. A SeparableConv2D $(3\times 3,256)$ [39] generates higher-dimensional features, followed by quantization-aware normalization and bounded activation:
$\widehat{F}=\min\left(1,\max\left(0,\gamma_{\text{spk}}\cdot\text{BN}(\text{SepConv}_{3\times 3,256}(F^{(4)}))+\beta_{\text{spk}}\right)\right)$ (7)
This mapping guarantees that all activations lie in $[0,1]$, which is both compatible with spike encoding and preserves information for subsequent inference.
To further optimize the channel-wise information flow, as shown in Fig. 5, a Squeeze-and-Excitation (SE) block [40] is applied, implementing a two-stage bottleneck and gating mechanism:
$\mathbf{s}=\sigma\left(W_{2}\,\delta\left(W_{1}\frac{1}{16}\sum_{i,j}\widehat{F}(i,j,:)\right)\right)$ (8)
where $W_{1}$ and $W_{2}$ are dense layers, $\delta$ is ReLU, and $\sigma$ is sigmoid. Each feature channel is then scaled by $s_{c}$, which adaptively modulates discriminative capacity and provides additional regularization for small, imbalanced datasets.
III-B3 Quantized Output Projection
The final block flattens the spike-compatible features $\widehat{F}$ to a vector $r\in\mathbb{R}^{4096}$ and applies a linear projection to produce the model output:
$\mathbf{y}=\mathbf{W}_{\text{cls}}\,r+\mathbf{b}_{\text{cls}}$ (9)
Here, $\mathbf{W}_{\text{cls}}\in\mathbb{R}^{7\times 4096}$, $\mathbf{b}_{\text{cls}}\in\mathbb{R}^{7}$. The output $\mathbf{y}$ is a $7$-dimensional, already quantized vector that can be seamlessly passed to the SNN converter for neuromorphic inference.
III-C CNN-to-SNN Conversion for Neuromorphic Deployment
The conversion of a trained convolutional neural network (CNN) into a spiking neural network (SNN) is a critical step for enabling event-driven inference on neuromorphic hardware. In this work, we utilize the Akida MetaTF toolkit [41, 15] to perform an automated and quantization-aware transformation of the CNN backbone described above into an SNN model suitable for direct deployment on the BrainChip Akida processor.
III-C1 Conversion Principles and Workflow
The Akida conversion [15] process follows a structured pipeline to ensure hardware compatibility and the preservation of model accuracy:
  • Operator Mapping: Each supported CNN layer (e.g., convolution, batch normalization, separable convolution, etc) is mapped to its spiking equivalent. For instance, ReLU activations are replaced by thresholding mechanisms that convert analog outputs to binary spike events. Layers incompatible with SNN operation (such as global average pooling or flatten) are substituted with spike-compatible alternatives, such as local pooling or spike-generating readout heads.
  • Quantization: All network weights and activations are quantized to a limited bit-width (8-bit), matching the precision constraints of the neuromorphic hardware. The quantization parameters are derived during the training and conversion phases to minimize information loss.
  • Temporal Spike Encoding: Continuous-valued activations from the CNN are converted into spike trains through rate coding or threshold-based event generation. The spike generation logic ensures that temporal information is preserved and the event-driven computation paradigm of SNNs is fully leveraged.
  • Resource and Constraint Adaptation: The MetaTF converter analyzes the input model to partition layers and neurons across the available hardware resources (neural cores, memory blocks) of the Akida chip, optimizing for parallelism, latency, and energy efficiency.
III-C2 Integration with Custom Backbone
Our proposed Quantization-Aware Network Architecture is specifically designed for seamless conversion. All intermediate activations are explicitly bounded and quantized, with network modules implemented in a form directly supported by Akida’s conversion pipeline. This ensures that no critical feature transformation is lost and that the functional mapping from input images to class predictions remains consistent between the CNN and its SNN counterpart.
III-C3 Conversion Output and Verification
Upon completion, the converter produces a deployable SNN model in Akida format. Model equivalence is empirically verified by comparing the output distributions of the original CNN and the converted SNN on a validation set. Any observed accuracy drop is mitigated through fine-tuning or incremental retraining on the SNN hardware, exploiting Akida’s support for on-chip learning.
III-D SNN Deployment, Optimization and Inference
The converted SNN model is deployed directly onto the BrainChip Akida neuromorphic processor for hardware-based inference. The deployment process comprises several steps: loading the SNN model into the hardware runtime environment, configuring input/output data streams, and initializing internal buffers and neuron state registers.
III-D1 Inference and Output Processing
Inference is performed in an event-driven manner, with input images encoded into spike trains and propagated through the SNN in a fully parallel fashion. For each test or validation sample, the model outputs spike counts or firing rates at the final output neurons, corresponding to the target classes. To robustly map temporal spike responses to class probabilities, we aggregate the spike counts $S_{c}(t)$ for each class $c$ within an integration window $T$, and apply a soft decision normalization:
$\hat{p}_{c}=\frac{\exp\left(\alpha\,\sum_{t=1}^{T}w_{t}S_{c}(t)\right)}{\sum_{k=1}^{C}\exp\left(\alpha\,\sum_{t=1}^{T}w_{t}S_{k}(t)\right)}$ (10)
where $w_{t}$ is an optional temporal weighting factor (e.g., $w_{t}=\exp(-\beta(T-t))$ for decaying integration), $\alpha$ is a scaling parameter, and $C$ is the number of output classes. Class prediction is then made as $\arg\max_{c}\hat{p}_{c}$.
III-D2 Parameter Calibration and On-Chip Optimization
After initial deployment, key runtime parameters—such as output spike thresholds, integration windows, and temporal pooling parameters—are empirically calibrated on a held-out validation set. To further optimize the class assignment and suppress spurious events, a threshold adaptation can be formulated as:
$\theta_{c}^{*}=\arg\min_{\theta}\left\{\sum_{i=1}^{N}\mathbb{I}\Big{[}\,y_{i}\neq\mathbb{I}\big{(}S_{c}^{(i)}>\theta\,\big{)}\Big{]}\right\}$ (11)
where $S_{c}^{(i)}$ is the total spike count for class $c$ on sample $i$, $y_{i}$ is the true label, and $\mathbb{I}[\cdot]$ is the indicator function. This allows data-driven threshold selection for each class. If performance deviation is observed relative to the original CNN, light on-chip fine-tuning is conducted using Akida’s incremental learning capability, adjusting only the last output layer to optimize for domain shift or quantization artifacts.
IV Experiment
This section presents the datasets, experimental setup, and evaluation protocols, followed by detailed quantitative analyses of classification performance, ablation studies, and efficiency metrics under various deployment conditions.
IV-A Datasets
We used two datasets: the public HAM10000 benchmark and a proprietary clinical dataset from Hospital Sultanah Bahiyah, Malaysia. Both sets reflect a range of lesion types and clinical diversity.
IV-A1 HAM10000 Dataset
The HAM10000 dataset [42] consists of 10,015 dermatoscopic RGB images labeled by expert dermatologists into seven categories: melanocytic nevus, melanoma, benign keratosis-like lesions, basal cell carcinoma, actinic keratosis/intraepithelial carcinoma, vascular lesions, and dermatofibroma. The images originate from diverse sources and exhibit significant class imbalance. For all experiments, we adopted a standard 70%/10%/20% train/validation/test split.
IV-A2 Hospital Sultanah Bahiyah Clinical Dataset
A proprietary clinical dataset was established in partnership with Hospital Sultanah Bahiyah, comprising 3,162 dermatoscopic images from 1,235 patients collected between June 2022 and February 2024. To ensure comparability with benchmark studies, lesion categories were selected to match the seven classes in HAM10000. Rare and unclassified lesions were excluded to maintain consistency in diagnostic labeling and facilitate joint evaluation. Each case was independently annotated by at least two board-certified dermatologists and histopathologically confirmed where possible. All images and metadata were anonymized following institutional ethical guidelines (approved protocol: NJG-2022-3233-CN). Dataset partitioning followed the same 70%/10%/20% train/validation/test split as HAM10000, stratified by disease category. This dataset provides a clinically diverse, real-world validation source for model generalization and robustness.
IV-B Experimental Setup
All model training, validation, and CNN-to-SNN conversion were conducted on a workstation with an Intel Core i9-12900K CPU, 128 GB RAM, and NVIDIA RTX 3090 GPU, running Ubuntu 22.04 LTS. The software environment included Python 3.9, CUDA 11.8, TensorFlow 2.10, and Akida MetaTF SDK v2.2.1. Neuromorphic inference was performed on a BrainChip Akida AKD1000 PCIe board installed in the same system, with deployment and testing managed via the Akida Python API and default board settings.
IV-C Analysis of Classification Results on HAM10000 Dataset
Table I reports the precision, recall, and accuracy of the proposed model for each class on the HAM10000 test set. The model achieves consistent performance across all lesion categories, including minority classes. Notably, the architecture maintains stable results under limited training data and on-chip incremental learning, enabling effective clinical adaptation in evolving or low-resource scenarios. With a Top-1 accuracy of 91.6% and macro F1 of 82.4%, the system demonstrates effective discrimination of both common and rare lesions.
TABLE I: Class-wise precision, recall, F1 score, and accuracy of QANA on the HAM10000 test set.
Class Precision Recall F1 Accuracy
Actinic keratoses 0.890 0.933 0.911 0.933
Basal cell carcinoma 0.890 0.901 0.896 0.901
Benign keratosis-like lesions 0.866 0.853 0.859 0.853
Dermatofibroma 0.925 0.976 0.950 0.976
Melanocytic nevi 0.887 0.817 0.851 0.817
Vascular lesions 0.949 0.966 0.957 0.966
Melanoma 0.956 0.933 0.944 0.933
Average 0.909 0.911 0.910 0.910
TABLE II: Performance Comparison of Converted SNN Models on HAM10000
Model (SNN, Akida) Top-1 Accuracy (%) Macro F1 (%)
ResNet-50 [9] 85.7 76.4
DenseNet-121 [32] 86.5 77.2
Inception-v4 [33] 85.9 76.9
EfficientNet-B4 [43] 87.3 78.1
MobileNet-v2 [30] 83.4 74.7
SENet-154 [28] 86.9 77.8
Xception [11] 85.5 76.2
Multi-Scale Attention [35] 87.0 78.0
CNN Ensemble [44] 88.1 78.9
AKIDANet [15] 83.2 73.6
Ours 91.6 82.4
IV-D Classification Performance on HAM10000
To comprehensively evaluate the effectiveness of our neuromorphic skin lesion classification system, we conducted a series of controlled experiments in which a selection of state-of-the-art convolutional neural network (CNN) architectures were converted to spiking neural networks (SNNs) using our quantization-aware pipeline and executed on the same Akida hardware platform. This strategy ensures a fair comparison of all approaches under identical hardware constraints and SNN deployment settings. The evaluated models include both canonical CNN baselines and advanced architectures commonly used in medical image analysis.
As shown in Table II, all evaluated CNN models show a decrease in accuracy and macro F1 score after conversion to SNN and deployment on neuromorphic hardware, with Top-1 accuracy values ranging from 83.2% to 88.1%. In contrast, our proposed model achieves a Top-1 accuracy of 91.6% and a macro F1 score of 82.4%, outperforming all baselines under the same SNN deployment conditions. This improvement demonstrates the benefit of our quantization-aware architecture and optimized network design for event-driven inference. The results confirm the practical utility of our method for real-time, resource-constrained medical applications, supporting efficient and accurate classification in portable diagnostic systems.
TABLE III: Per-image inference latency and energy consumption of all models on the HAM10000 test set. CNN baselines are measured on an NVIDIA RTX 3090 (GPU) and Intel Xeon Gold 6226R (CPU); SNNs are measured on BrainChip Akida AKD1000. All values are averaged over 10,000 images. The rightmost columns show the percentage reduction achieved by SNNs on Akida compared to the corresponding CNN (GPU) version.
Model CNN (GPU) CNN (CPU) SNN (Akida) Relative Reduction(%)
Latency (ms) Energy (mJ) Latency (ms) Energy (mJ) Latency (ms) Energy (mJ) Latency Energy
ResNet-50 [9] 12.1 175.2 57.9 923.3 2.8 3.3 76.9 98.1
DenseNet-121 [32] 14.7 199.6 68.4 1075.2 3.1 3.5 78.9 98.2
Inception-v4 [33] 16.8 218.5 82.1 1237.5 3.5 4.1 79.2 98.1
EfficientNet-B4 [43] 18.9 242.1 93.6 1345.6 4.0 4.6 78.8 98.1
MobileNet-v2 [30] 6.9 97.5 68.2 1082.1 2.2 2.6 68.1 97.3
SENet-154 [28] 17.3 236.8 89.4 1312.4 3.8 4.4 78.0 98.1
Xception [11] 10.7 151.2 51.5 863.7 2.6 3.1 75.7 97.9
Multi-Scale Attention [35] 19.8 251.7 97.7 1391.5 4.2 5.0 78.8 98.0
CNN Ensemble [44] 36.1 450.5 157.2 2177.1 7.5 8.6 79.2 98.1
QANA (Ours) 27.6 163.1 83.9 841.5 1.5 1.7 94.6 98.6
TABLE IV: Ablation study of core modules in our model on the HAM10000 test set. Metrics are reported as percentages (%). Each row shows the incremental addition of modules.
Configuration Ghost Block ECA SE Quant. Head SMOTE Inc. Learn Accuracy Recall Precision F1 Score AUC-ROC
Baseline 74.1 71.4 71.9 71.6 77.3
+ Ghost Block 72.3 70.2 70.0 70.6 70.9
+ ECA 88.7 85.8 87.2 86.5 90.7
+ SE 89.8 87.7 88.1 87.8 91.5
+ Augmentation 90.4 88.1 89.1 88.6 92.1
+ SMOTE 91.0 89.2 90.0 89.6 92.7
+ Incremental Learning 91.6 90.7 91.2 90.9 93.4
IV-E Inference Speed and Energy Consumption
We quantitatively evaluated the inference latency and energy consumption of our neuromorphic model in comparison with both conventional CNN baselines and other SNN-converted architectures. All SNN models were deployed on the Akida AKD1000 PCIe board, while CNN baselines were tested on both NVIDIA-RTX-3090-GPU and an Intel Xeon CPU. For each model, the reported inference latency corresponds to the average per-image processing time over 10,000 test samples. Energy consumption per image was measured as the mean device power during inference multiplied by the average inference time, using on-board power monitoring for Akida and NVIDIA-smi for GPU and Intel RAPL for CPU.
To ensure a fair and hardware-consistent comparison, all tested CNN architectures were converted to SNNs using our pipeline before deployment on Akida. Table III presents a comprehensive summary of inference latency and energy consumption across all evaluated models and platforms, as well as the relative reduction of these metrics for SNNs on Akida compared to their CNN GPU and CPU implementations. Our neuromorphic model achieves the lowest inference latency and energy consumption of all evaluated architectures. When deployed as an SNN on the Akida platform, it completes classification in 1.5 ms per image and requires just 1.7 mJ, representing an 94.6% reduction in latency and over 99.0% reduction in energy compared to the equivalent CNN on GPU surpassing all other state-of-the-art CNN-to-SNN conversion baselines.
TABLE V: Performance comparison of converted SNN models on the Clinical Dataset.
Model (SNN, Akida) Top-1 Accuracy (%) Macro F1 (%)
ResNet-50 [9] 84.6 75.3
DenseNet-121 [32] 85.7 76.2
Inception-v4 [33] 85.2 75.8
EfficientNet-B4 [43] 86.3 77.0
MobileNet-v2 [30] 82.8 73.7
SENet-154 [28] 85.4 76.6
Xception [11] 84.2 74.8
Multi-Scale Attention [35] 86.5 77.2
CNN Ensemble [44] 87.6 78.1
AKIDANet [15] 81.9 71.5
Ours 90.8 81.7
IV-F Ablation Study of Model Components
Table IV reports the ablation results on the HAM10000 test set, illustrating the contribution of each core module to the overall performance. Modules were incrementally enabled to measure their isolated and cumulative effects on classification metrics, including accuracy, recall, precision, F1 score, and AUC-ROC (%). The results demonstrate that each module provides consistent improvements, with the complete model achieving the highest accuracy and F1 score.
IV-G Classification Performance on Clinical Dataset
We evaluated the proposed neuromorphic skin lesion classification system and several representative CNN architectures, all converted to SNNs and deployed on the Akida hardware platform. Table V presents the Top-1 accuracy and macro F1 score for each model under the same deployment conditions. Our model achieves the highest accuracy and macro F1 score among all tested methods, confirming its effectiveness and robustness for neuromorphic inference in clinical scenarios.
V Conclusion
In this paper, we proposed QANA, a quantization-aware neuromorphic framework for skin lesion classification on edge devices. Extensive experiments on the large-scale HAM10000 benchmark and a real-world clinical dataset show that QANA achieves state-of-the-art accuracy (91.6% Top-1, 82.4% macro F1 on HAM10000; 90.8%/81.7% on the clinical set) while enabling real-time and energy-efficient inference on the BrainChip Akida platform (1.5 ms latency, 1.7 mJ per image). These results demonstrate that QANA is highly effective for portable medical analysis and AI deployment in dermatology under limited computing resources.
VI Acknowledgement
This research was supported by the National Natural Science Foundation of China (Nos. 62172336 and 62032018). The authors gratefully acknowledge BrainChip Holdings Ltd. for providing technical support and the Akida AKD1000 hardware platform, whose powerful neuromorphic computing capabilities enabled strong performance of the SNN model. The authors also extend their appreciation to Dr. Atif Mansoor, Dr. Bo Miao and their teams for their preliminary contributions to this research.